A hairdresser has 13 small and 14 large cans of hairspray, giving a total of 9 L of hairspray. At this time last year she had 4 small and 12 large cans, totaling 6 L hairspray. How much spray is in each size can?
Answer by richwmiller(17219)

13s+14l=9,
4s+12l=6
l = 21/50=.42 liter, s = 6/25=12/50=.24 Liter